FIND THE GENERAL SOLUTION.Find the solution of two dimensional Laplaces equation: 227 227 + дх2 = 0 ay2 Use separation of variables and Fourier series
Answer #1

דל 2²7 de + Jyr=o. let T (1,4)= X(a). Y14) 2T да 37 = xY A - XY @ AT sy 37 exy (3) syn = xy? ☺.☺ in xy + y X=0 osph u xCose 2 when M=0. X=0 Y:0 X = 62+2 Y = czy + c px T=xY To (ca 14) Cusy the 16 case 3 when M20 - ²M x+ix zo os X1 + x² | 9-please provide feedback:)